OpenStudy (anonymous):

What is t in tan(t)=2 ? Im using a TI82 but it gives me a weirds numbers.

OpenStudy (amistre64):

i know even and odd numbers; whats a wierd number?

OpenStudy (anonymous):

t = @12*\[\pi\]-tan^-1(1/2) + pi/2

OpenStudy (anonymous):

\[e^{-\pi i \text{weird}}\]

OpenStudy (anonymous):

It doesnt make sense

OpenStudy (amistre64):

have you tried replacing the batteries?

OpenStudy (phi):

are you in radian mode?

OpenStudy (anonymous):

Yes Im in radian mode

OpenStudy (amistre64):

if the radiation badge has turned colors, its time to get a new ti82

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=arctan%282%29

OpenStudy (anonymous):

My TI89 (It's 89) is new

OpenStudy (anonymous):

Can anyone solve t for me?

OpenStudy (phi):

then you should get something like 1.107 rads

OpenStudy (zarkon):

you did solve for t... you have the answer

OpenStudy (anonymous):

No I know it's not the answer....

OpenStudy (zarkon):

lol...it is

OpenStudy (zarkon):

it's just not a form you like

OpenStudy (amistre64):

\[tan^{-1}(x) =x-\frac{x^3}{3}+\frac{x^5}{5}-\frac{x^7}{7}+\frac{x^9}{9}-\frac{x^{11}}{11}+\frac{x^{13}}{13}... \] just plug in x=2

OpenStudy (amistre64):

thats the way we had to do it back in the monastary

OpenStudy (anonymous):

I get it thanks

OpenStudy (anonymous):

satellite73 i need your help on the last problem you helped me with.

OpenStudy (zarkon):

unless you specify the domain you are working over you will have an infinite number of solutions...just as your calculator indicated

OpenStudy (anonymous):

It's part of this question Find the maximum and minimum values of f(x,y) = x+2y on the disk x^2+y^2 ≤1 I have done this for now: f_1(x,y) = 1 f_2(x,y) = 2 x=cos(t) and y=sin(t) I have that g(t) = x(t) + 2*y(t) --> g(t) = cost(t) + 2*sin(t) g'(t) = 0 = 2*cost-sin(t) Then I can see that: 2cos(t)/cos(t) -sin(t)/cos(t) = 0/cos(t) --> tan(t) = 2 From this point I have no idea what to do.

OpenStudy (zarkon):

use lagrange multipliers

OpenStudy (anonymous):

That is the given solution....I just cant follow what they are doing

OpenStudy (zarkon):

or...just compute \[\cos(\tan^{-1}(2))+2\sin(\tan^{-1}(2))=\sqrt{5}\]

OpenStudy (anonymous):

Thank you a lot you just helped me finish this assignment! Greetings from Denmark

OpenStudy (phi):

In case anyone is interested, though I'm not sure it sheds much light.